### v1

# Let and let be an open subset of Let and let be a family of smooth vector fields on which span the tangent space at every point of Let be positive integers. Let be the usual (Euclidean) norm on 5.1. Metric associated to a family of vector fields Let be an -dimensional vector space, and equip with the family of dilations defined by

Let and let be an open subset of Let and let be a family of smooth vector fields on which span the tangent space at every point of Let be positive integers. Let be the usual (Euclidean) norm on
5.1. Metric associated to a family of vector fields
Let be an -dimensional vector space, and equip with the family of dilations defined by
That is,
Fix a norm on which is homogeneous with respect to i.e. for all
Let Note that
We will write for the ball in with respect to the homogeneous norm
Since the Lebesgue measure of is proportional to
Let denote the linear map from to the Lie algebra of vector fields on given by
For any there exists such that the map defined by
exists and is a submersion. Here, is the exponential of vector fields (i.e., the flow at time ). We can see that exists for sufficiently small because is smooth. It is also clear that the derivative of at is surjective since spans the tangent space at Hence, exists and is a submersion for sufficiently small For further details, see [Goo76, Page 39].
Since is a submersion, the pre-image of a point is a (possibly empty) -dimensional submanifold of Note that whether or not is empty will depend on the choice of
Let be the Jacobian matrix of since is a submersion, is a -matrix of full rank. Define
Here, is the matrix transpose.
The co-area formula asserts that
Here is the -dimensional Hausdorff measure on the submanifold
See [Fed69, Theorem 3.2.12]. We will frequently refer to this measure and so we introduce the following notation.
Definition 5.1.
Let Let denote the following measure on
With this notation, the coarea formula for asserts that
We record a few essential features of
Lemma 5.2.
Let We have
and on
Proof.
By definition, if and only if and Since we have and hence if we have
This proves that if and then
By symmetry, we conclude the converse, and this proves the first claim.
The fact that on follows from (5.1).
β
Using we can also define the following function on
Definition 5.3.
Let If define
Otherwise,
Equivalently, if and if there exists such that Observe that is precisely the function denoted in [NSW85, Section 4]. The same metric was used by Street, see [Str23, Section 1.3].
We want to use as a quasi-metric on As defined, may fail to be a quasi-metric for the obvious reason that may be infinite. A more subtle problem is that is not necessarily symmetric: it might be that but
We deal with this issue by assuming that is as small as necessary. Theorem 2 of [NSW85] implies the following:
Theorem 5.4.
Every point has a neighbourhood such that for all and Moreover, for all
Henceforth, replacing with if necessary, we will make the following assumption:
Assumption 1.
Assume that is small enough so that is finite and symmetric on and also so that
Replacing by we assume without loss of generality that is constant for
We also assume that has compact closure, in order that all of the derivatives of with respect to are uniformly bounded over
Ultimately there will be no harm in making this assumption, since our final estimates will take place over compact subsets of which can be covered by a finite number of sets satisfying Assumption 1.
Using this new assumption, is a non-empty submanifold of for every
Now that is guaranteed to be finite and symmetric on we can use the following theorem.
Theorem 5.5 (Nagel-Stein-Wainger).
[NSW85, Section 4]
As given in Definition 5.3, defines a quasi-metric on
which is uniformly comparable to the Euclidean metric in the sense that there exist constants such that
If denotes the Lebesgue measure on then is a doubling quasi-metric measure space.
We will need some estimates on integrals with respect to the measures
Lemma 5.6.
Let and
There exists a constant such that
and also
If then
Proof.
If then (5.2), implies that so
It follows that for small the two claims in (1) are equivalent. The same follows for general by increasing the constant
By the coarea formula (5.1), if then
By increasing if necessary, this still holds for
This proves (1).
For (2), note that by definition if we have and so
β
5.2. Graded basis
Let be as above. In particular, we continue to enforce Assumption 1.
Let be an open subset of which is invariant under the -action defined by
The -tuple is said to a graded basis for the filtration defined by in the terminology of [AMY22, Definition 2.1].
Definition 5.7.
Let denote the projection map
A function on is said to be -properly supported if is proper on the support of That is, if for any compact subset of is compact.
Definition 5.9.
Let be -properly supported.
Assume that there exists such that
Assume that if and then is supported in
For fixed and define
That is, is a function on such that for any we have
Let
Our goal is to show that in Definition 5.9 obeys all of the conditions of Definition 3.1, with where is the Lebesgue measure. Conditions (I), (II) and (IV) are easy consequences of the Lemma 5.6, while Condition III will be more challenging.
Note that the condition that there exists such that
is equivalent to
Indeed, since is smooth there exists an -properly supported smooth function such that
and as
Lemma 5.10.
Let and be as in Definition 5.9. There exists a constant such that if
then That is, satisfies Condition (I) of Definition 3.1 with constant
Proof.
By assumption, we have
Hence, the result follows from Lemma 5.6.(2).
β
Lemma 5.11.
Let and be as in Definition 5.9. There exists a constant , independent of such that
That is, satisfies Condition (II) of Definition 3.1, with constant
Proof.
Exactly as in Lemma 5.10, we have
The result follows from Lemma 5.6.(1).
β
Lemma 5.12.
Let and be as in Definition 5.9. There exists a constant independent of such that
That is, satisfies Condition (IV) of Definition 3.1, with constant
Proof.
By the coarea formula (5.1) and the assumption on we have
β
5.3. Level sets of submersions
The final and most difficult condition to verify from Definition 3.1 is Condition (III). In order to achieve this, we need to understand how the measure varies with This can be done using classical vector calculus.
To simplify the notation, in this subsection we consider a general submersion
where recall that and and is equipped with a norm The fact that is equipped with a family of anisotropic dilations will not be important in this subsection. Ultimately we will consider which is only defined on an open subset of but for brevity we work here with defined on the entire space. Since our reasoning is local this is no loss of generality.
Given let denote the measure on given by
The coarea formula asserts that
In this subsection we will derive a formula for the difference
The intent will be to analyse the difference where is as in Definition 5.9.
The idea is to define a flow in which transports to
Let denote the vector field on given by
Since is a submersion and the vector field is nowhere vanishing. Let be the flow of this vector field. That is, or more specifically
solves the differential equation
In principle, the flow may only exist for a time
If it exists, the map transports to One way to see this is to define
The chain rule gives
Hence, if then for all
It follows that defines a smooth map from to In particular, at we have a smooth map
Similar reasoning shows that is a smooth inverse for so is a diffeomorphism.
The next step is to determine how changes with Again let We have
Here, is the pullback of to and is the determinant.
To see (5.5), let and By (5.3), we have
By a change of variables
Applying (5.3) again gives
Since and are arbitrary, (5.5) follows.
Using (5.5), we have
Therefore
Recall that where is the vector field (5.4). The Liouville-Ostrogradsky formula says that Jacobian determinant of obeys the differential equation
Since we have and hence
Thus,
With this preparation, we can prove the main result of this subsection.
Proposition 5.13.
Let be a submersion, let
There exists and such that for all ,
and we have
Proof.
Let be the vector field in (5.4).
Since is smooth and depends linearly on for every and and there is a constant such that
Let By the Picard-LindelΓΆf theorem, if
then for the flow exists and moreover
This implies that
Similarly,
Since
we have
Since this implies
Therefore,
Similarly,
Let Note that is supported in since
Now we apply (5.3), resulting in
By (5.8), (5.7)
and the numerical inequality
we have
Since we conclude the result with constant
β

### v3

@@
-Let and let be an open subset of Let and let be a family of smooth vector fields on which span the tangent space at every point of Let be positive integers.
-Let be the usual (Euclidean) norm on 5.1.
-Metric associated to a family of vector fields Let be an -dimensional vector space, and equip with the family of dilations defined by
+estimates in the Androulidakis-Mohsen-Yuncken calculus
+Abstract.
+We prove that order zero operators in the pseudodifferential calculus associated to a filtration defined by Androulidakis, Mohsen and Yuncken are bounded on spaces for
+1. Introduction
+Recently, Androulidakis, Mohsen and Yuncken introduced a pseudodifferential calculus adapted to an arbitrary filtration of a manifold [AMY22].
+Using this calculus, they resolved the long-standing Helffer-Nourrigat conjecture characterising maximally hypoelliptic operators in terms of representation theory.
+We describe their setting as follows.
+Let be open, and let be smooth vector fields on satisfying the HΓΆrmander bracket condition on That is, the Lie algebra of vector fields on generated by spans the tangent space of at every point.
+Equip each vector with a weight This data induces a filtered foliation on which is a filtration of the algebra of vector fields by locally finitely generated -submodules
+Here, is the -linear span of all the commutators of the form
+Let The -order of a differential operator is the least integer such that can be written in the form
+where is the operator of pointwise multiplication by a smooth function on and
+Let denote the space of differential operators of -order at most
+The algebra of differential operators on is a filtered algebra in the sense that
+In [AMY22], the authors constructed a filtered algebra extending which has most of the desirable properties of a pseudodifferential calculus.
+The construction was based on an βapproximate homogeneityβ condition first described by Debord and Skandalis [DS14], and later used as the basis of pseudodifferential calculus by van Erp and Yuncken [vEY19].
+One of the desirable properties of a pseudodifferential calculus is that operators of order zero should act as bounded operators on the space of square integrable functions, at least locally.
+This property for is proved in [AMY22, Theorem 3.31(b)].
+The proof there followed the classic strategy of first showing that an operator with positive principal symbol has an approximate square root.
+Since symbols in this calculus are not functions but instead elements of some -algebra, showing that the desired approximate square root belongs to the calculus is not straightforward.
+It is desirable to have a more self-contained argument.
+This is what we supply here.
+An off-shoot of this effort is that we also show that operators in are also bounded on spaces for in a suitable local sense.
+1.1. Acknowledgements
+The key idea in the proof of the main theorem, and arguably the main novelty of this paper, is the decomposition of kernels according to the -action described in the proof of Theorem 6.5.
+The observation that this decomposition is almost orthogonal is due to Omar Mohsen, who kindly shared it with us.
+Thanks are also due to RaphaΓ«l Ponge for helpful comments and suggestions.
+2. Main results
+Let be a -manifold, and let be a filtration of the Lie algebra of vector fields on by -submodules.
+Let be the pseudodifferential operators of order on in the sense of Androulidakis, Mohsen and Yuncken.
+We will recall the definition of this calculus is recalled in Section 5 below.
+A pseudodifferential operator is associated to a one-parameter family with This extension is unique up to the addition of a smoothly varying family of smoothing operators.
+Given the -space can be defined invariantly as a space of -densities on Alternatively, we can choose a fixed nonzero density on and identify with a space of functions.
+Since our results are local, this distinction is not important.
+Given we write for the operator of pointwise multiplication by on We say that an operator is locally bounded on if is bounded from to for all
+Theorem 2.1.
+Let and let If then is locally bounded on Moreover, the associated family is uniformly continuous as in the sense that if then
+If then the family is uniformly continuous as on for any and moreover is locally compact in the sense that if then is compact on for
+Remark 2.2.
+Concerning the proof, we make a few remarks.
+β’
+With there is no loss of generality in considering since the closure of the calculus under composition and adjoint implies that if then For the proof needs to allow general
+β’
+The proof is based on an almost-orthogonal decomposition, and the operator is decomposed into a sum The proof of Theorem 2.1 shows that this sum is unconditional in
+β’
+In fact in the theorem obeys a weak-type estimate, uniformly as The Marcinkiewicz interpolation theorem implies that the norm of from to is bounded by a constant multiple of
+It should be noted that Theorem 2.1 is not new.
+A very general class of operators was proved by Street to be bounded on -spaces in [Str14].
+Those operators are defined as having symbol functions that obey standard estimates [Str14, Definition 4.1.1], and the operators in the Androulidakis-Mohsen-Yuncken calculus have kernels that obey similar estimates, this follows from [AMY22, Proposition 3.4] by an argument similar to [vEY19, Corollary 45].
+Therefore, are order zero operators in Streetβs sense and Theorem 2.1 follows from [Str23, Theorem 5.10.1].
+In fact, Street also proves the boundedness on spaces of so-called multiparameter pseudodifferential operators.
+Since these are typically not of weak-type (for example, consider a tensor product of two pseudodifferential operators), the proof of Theorem 2.1 given here does not extend to that case.
+Nevertheless, the main reason for this paper is to give a proof of Theorem 2.1 within the context of the Androulidakis-Mohsen-Yuncken calculus.
+An important special case occurs when the filtration arises as the sections of a filtration of by vector bundles.
+This was the case considered in [vEY19].
+The boundedness of order zero operators on spaces in this case follows from Theorem 2.1, but in this case similar statements have long been known, see [Goo76, Page 103].
+The proof of Theorem 2.1 is mostly independent of the results in [AMY22].
+The only results used from there are that the calculus is closed under composition and the adjoint.
+The -boundedness of order zero operators was proved in [AMY22], but here we obtain the same result by other methods.
+Associated to the calculus is a scale of Hilbert-Sobolev spaces adapted to the filtration [AMY22, Section 3.5].
+Having the boundedness of operators of order zero -spaces, we can develop a similar theory for -Sobolev spaces.
+We supply some aspects of this theory in Section 8, including a Sobolev embedding theorem and in Section 9 deduce some results concerning maximal subellipticity in for HΓΆrmander operators.
+3. Abstract kernel estimates
+In this section and the next we will work in the setting of abstract harmonic analysis, similar to [CW71, DH09].
+The style of argument will be familiar to experts in harmonic analysis and similar arguments have appeared elsewhere, see in particular Goodman [Goo76].
+Let be a quasi-metric doubling measure space. That is, there exists a constant such that
+and there exists a constant such that
+where denotes the ball about of radius with respect to
+Ultimately, will be a subset of Euclidean space, will be the Lebesgue measure and will be a quasi-metric defined by Nagel, Stein and Wainger [NSW85], but for this section and the next we work in an abstract setting.
+We want to avoid making the assumption that there exist constants such that
+This assumption was made in [Goo76], but it is unclear whether it holds in our situation.
+We will consider singular integral operators with kernels given formally as infinite linear combinations We will require in different places various conditions on the individual terms listed as Conditions (I)-(IV), defined as follows.
+Definition 3.1.
+Let be a sequence of locally integrable functions on We consider the following conditions on the sequence :
+There is such that for all is supported in the set of where
+There is such that for all we have
+There are constants and and a sequence of functions on such that if then
+and
+There is such that
+Given a finitely supported sequence denote
+Also let be the sequence
+Some initial remarks on the definition:
+β’
+(I) is indifferent to the order of and so that if it applies to the functions then it also applies to with the same constant.
+β’
+All four conditions apply with the same constants if we replace by or where
+β’
+If satisfies (I), then for and so the series defining makes sense for arbitrary sequences when
+β’
+The constant in Condition (III) is denoted differently to because the size of the constant has a different effect on the strength of the condition.
+Whereas a smaller value of for implies a stronger condition, a smaller value of implies a weaker condition.
+We start with the following relatively easy consequence of the first two conditions in Definition 3.1
+The following shows that the sums from Definition 3.1 obey the estimates required of a CalderΓ²n-Zygmund kernel as in [CW71, Chapter III, Theorem (2.4), p.74].
+Lemma 3.3.
+Proof.
+Note that in region of integration, is only evaluated when Since is non-zero only when the sum
+is finite for every
+By definition,
+We split the sum into parts where and where In the region of integration, we have and hence by the quasi-triangle inequality
+Therefore If then we have both and greater than Hence,
+Thus
+Applying Condition (III) to the first summand and Condition (II) to the second summand gives
+By the numerical inequalities
+and
+we conclude
+where depends only on on and on
+β
+In order to obtain boundedness for we need to verify an almost-orthogonality condition on the kernels and also their adjoints.
+Theorem 3.4.
+Let and both satisfy all of the conditions of Definition 3.1. Assume without loss of generality that for and Then
+where depends only on on on and on
+Proof.
+Let
+Note that
+We will show separately that
+and
+The combination of (3.2), (3.3) and (3.4) implies the result.
+First, to show (3.3), we write as a sum as follows:
+Exchanging the order of integration in the first term gives
+Applying (II) for and (IV) for to the second summand gives
+Since it follows that
+We therefore concentrate on bounding the first summand above. Condition (III) implies that
+Now we prove (3.4). Write
+Exchanging the order of integration we obtain
+By (II) and (IV) applied to the second term is bounded above by We therefore concentrate on the first summand.
+Since within the support of we have Applying (III) to and (3.1) to we obtain
+This completes the proof of (3.3).
+β
+4. Abstract Operator estimates
+In this section we continue in the same setting as Section 3: is a quasi-metric space, and is a doubling measure on with respect to
+Let be a kernel satisfying all of the conditions of Definition 3.1, and assume that the same holds for Without loss of generality, we assume throughout that for and
+Given a locally integrable function on denote for the integral operator with kernel That is,
+Theorem 4.1.
+Proof.
+Assume without loss of generality that We have
+Note that
+By Theorem 3.4,
+Dually, we have
+By complex interpolation, it follows that
+Exchanging the roles of and gives
+That is,
+The -boundedness of the sum follows from this estimate and Cotlar-Knapp-Stein almost-orthogonality. β
+Corollary 4.2.
+Proof.
+Corollary 4.3.
+Let and satisfy all of the conditions of Definition 3.1 and let be a finitely supported sequence. Then is bounded on for all and there exists a constant depending only on the quantities
+Moreover,
+Proof.
+It is also worth mentioning the following much easier estimate.
+Proof.
+By subadditivity, it is enough to prove the result for a single The assumption (II) asserts the result for and since the same holds for we also have the result for The general case follows from interpolation.
+β
+5. Kernels derived from graded bases
 Let and let be an open subset of Let and let be a family of smooth vector fields on which span the tangent space at every point of Let be positive integers. Let be the usual (Euclidean) norm on
@@
 That is,
-Fix a norm on which is homogeneous with respect to i.e. for all
-Let Note that
-We will write for the ball in with respect to the homogeneous norm
-Since the Lebesgue measure of is proportional to
+Fix a norm on which is homogeneous with respect to i.e. for all Let Note that We will write for the ball in with respect to the homogeneous norm Since the Lebesgue measure of is proportional to
 Let denote the linear map from to the Lie algebra of vector fields on given by
@@
 It is also clear that the derivative of at is surjective since spans the tangent space at Hence, exists and is a submersion for sufficiently small For further details, see [Goo76, Page 39].
-Since is a submersion, the pre-image of a point is a (possibly empty) -dimensional submanifold of Note that whether or not is empty will depend on the choice of
-Let be the Jacobian matrix of since is a submersion, is a -matrix of full rank. Define
-Here, is the matrix transpose.
-The co-area formula asserts that
-Here is the -dimensional Hausdorff measure on the submanifold
-See [Fed69, Theorem 3.2.12]. We will frequently refer to this measure and so we introduce the following notation.
+Since is a submersion, the pre-image of a point is a (possibly empty) -dimensional submanifold of Note that whether or not is empty will depend on the choice of Let be the Jacobian matrix of since is a submersion, is a -matrix of full rank.
+Define
+Here, is the matrix transpose. The co-area formula asserts that
+Here is the -dimensional Hausdorff measure on the submanifold See [Fed69, Theorem 3.2.12]. We will frequently refer to this measure and so we introduce the following notation.
 Definition 5.1.
@@
 Proof.
-By definition, if and only if and Since we have and hence if we have
-This proves that if and then
-By symmetry, we conclude the converse, and this proves the first claim.
-The fact that on follows from (5.1).
-β
+By definition, if and only if and Since we have and hence if we have This proves that if and then By symmetry, we conclude the converse, and this proves the first claim.
+The fact that on follows from (5.1). β
 Using we can also define the following function on
@@
 Theorem 5.5 (Nagel-Stein-Wainger).
-[NSW85, Section 4]
-As given in Definition 5.3, defines a quasi-metric on
-which is uniformly comparable to the Euclidean metric in the sense that there exist constants such that
-If denotes the Lebesgue measure on then is a doubling quasi-metric measure space.
 We will need some estimates on integrals with respect to the measures
@@
 By the coarea formula (5.1), if then
-By increasing if necessary, this still holds for
-This proves (1).
-For (2), note that by definition if we have and so
-β
+By increasing if necessary, this still holds for This proves (1).
+For (2), note that by definition if we have and so β
 5.2. Graded basis
 Let be as above. In particular, we continue to enforce Assumption 1.
-Let be an open subset of which is invariant under the -action defined by
+Let be an open subset of containing which is invariant under the -action defined by
 The -tuple is said to a graded basis for the filtration defined by in the terminology of [AMY22, Definition 2.1].
@@
 A function on is said to be -properly supported if is proper on the support of That is, if for any compact subset of is compact.
+Remark 5.8.
+If then saying that is -properly supported is simply saying that it is compactly supported in the -variable. For example, is -properly supported precisely when we can identify with a function
 Definition 5.9.
-Let be -properly supported.
-Assume that there exists such that
-Assume that if and then is supported in
+Let be -properly supported. Assume that if and then is supported in
 For fixed and define
@@
 Conditions (I), (II) and (IV) are easy consequences of the Lemma 5.6, while Condition III will be more challenging.
-Note that the condition that there exists such that
-is equivalent to
+Remark 5.10.
+An additional condition that we may place on , besides Definition 5.9, is that there exists such that
+This is equivalent to
 Indeed, since is smooth there exists an -properly supported smooth function such that
@@
 Lemma 5.11.
-Let and be as in Definition 5.9. There exists a constant such that if
-then That is, satisfies Condition (I) of Definition 3.1 with constant
-Proof.
-By assumption, we have
-Hence, the result follows from Lemma 5.6.(2).
-β
 Lemma 5.12.
-Let and be as in Definition 5.9. There exists a constant , independent of such that
-That is, satisfies Condition (II) of Definition 3.1, with constant
-Proof.
-Exactly as in Lemma 5.10, we have
-The result follows from Lemma 5.6.(1).
-β
 Lemma 5.13.
-Let and be as in Definition 5.9. There exists a constant independent of such that
-That is, satisfies Condition (IV) of Definition 3.1, with constant
-Proof.
-By the coarea formula (5.1) and the assumption on we have
-β
 5.3. Level sets of submersions
@@
 The intent will be to analyse the difference where is as in Definition 5.9.
-The idea is to define a flow in which transports to
-Let denote the vector field on given by
-Since is a submersion and the vector field is nowhere vanishing. Let be the flow of this vector field. That is, or more specifically
-solves the differential equation
+The idea is to define a flow in which transports to Let denote the vector field on given by
+Since is a submersion and the vector field is nowhere vanishing. Let be the flow of this vector field. That is, or more specifically solves the differential equation
 In principle, the flow may only exist for a time
@@
 The chain rule gives
-Hence, if then for all
-It follows that defines a smooth map from to In particular, at we have a smooth map
+Hence, if then for all It follows that defines a smooth map from to In particular, at we have a smooth map
 Similar reasoning shows that is a smooth inverse for so is a diffeomorphism.
 The next step is to determine how changes with Again let We have
-Here, is the pullback of to and is the determinant.
-To see (5.5), let and By (5.3), we have
+Here, is the pullback of to and is the determinant. To see (5.5), let and By (5.3), we have
 By a change of variables
@@
 Since and are arbitrary, (5.5) follows.
-Using (5.5), we have
-Therefore
-Recall that where is the vector field (5.4). The Liouville-Ostrogradsky formula says that Jacobian determinant of obeys the differential equation
+Recall that where is the vector field (5.4).
+The Liouville-Ostrogradsky formula111This formula appears to be folklore.
+A proof is sketched as an exercise on Page 86 of [Per01].
+says that Jacobian determinant of obeys the differential equation
 Since we have and hence
@@
 Proposition 5.14.
-Let be a submersion, let
-There exists and such that for all ,
-and we have
+Let be a submersion, let There exists and such that for all , and we have
 Proof.
-Let be the vector field in (5.4).
-Since is smooth and depends linearly on for every and and there is a constant such that
+Let be the vector field in (5.4). Since is smooth and depends linearly on for every and and there is a constant such that
 Let By the Picard-LindelΓΆf theorem, if
@@
 Similarly,
-Let Note that is supported in since
-Now we apply (5.3), resulting in
-By (5.8), (5.7)
-and the numerical inequality
-we have
-Since we conclude the result with constant
+5.4. The proof of (III) for
+Theorem 5.15.
+Proof.
+For brevity, let
+Recall that is supported in and since and there is a constant such that
+By the definition of
+Now we apply Proposition 5.14 with and and If is chosen as in Proposition 5.14, then for we have
+Since depends smoothly on and has been chosen as in Assumption 1, the constants and can be chosen uniform over By (5.2), we can replace the Euclidean distance by Hence, for we have
+By (5.2), this still holds if for sufficiently small and this completes the proof.
 β
+Corollary 5.16.
+Proof.
+The results of this section are summarised in the following theorem:
+6. The Androulidakis-Mohsen-Yuncken calculus
+In this section we recall the definition of the Androulidakis-Mohsen-Yuncken calculus, and prove Theorem 2.1.
+Let be as in the previous section.
+That is, is open, is an -dimensional vector space equipped with dilations is an open subset of containing and invariant under the action
+and is a linear map from to vector fields on
+As previously, is defined as Note that is surjective.
+We consider distributions on which are -fibred and -proper. The set of such distributions is written Equivalently, is a linear map
+which is -linear, where is a -module in the obvious way.
+We will consider distributions as generalised functions rather than generalised densities. That is, if is a smooth function on then determines an -fibred distribution by
+where is the Lebesgue measure on This differs from the convention in [vEY19], where the measure was incorporated into the distribution.
+Given an -fibred distribution the pushforward is defined on functions by and on -fibred distributions by
+Remark 6.1.
+If then can be described slightly more simply. We have a continuous identification
+That is, if is smooth in and compactly supported in
+An -fibred distribution defines a family of linear operators from to by
+This was denoted in [AMY22].
+Note that a priori this formula may not make sense for all and because the exponential may not be defined.
+For this reason we assume that and restrict attention to a set satisfying Assumption 1.
+By the co-area formula, we have
+That is, with respect to the Lebesgue measure on has integral kernel
+where is the same as in Definition 5.1. Note that if is smooth away from formula makes sense for every
+The following is an adaptation of [AMY22, Definition 3.3].
+Definition 6.2.
+Let be as above. For the set is the subspace of -fibred distributions such that for all we have
+This is an βadaptationβ of the definition in [AMY22] rather than a direct translation because the operators there were considered to act on half-densities, whereas here we consider operators acting on functions.
+This causes a shift by in the exponent of but this does not change the class of operators.
+Definition 6.3.
+For the set of pseudodifferential operators associated to the filtration defined by with weights is the set of operators of the form where
+and
+Given an associated family is any choice defined by (6.1) for which and
+For denotes the -space with respect to the Lebesgue measure on that is
+As usual this should be replaced with the space of essentially bounded functions on when
+Definition 6.4.
+For , we say that an operator is locally bounded on if for all smooth compactly supported functions there is a constant such that
+Similarly, say that is locally compact if is compact on for all We say that a family of operators is uniformly continuous as if the constant can be chosen bounded above as That is, if there exists a constant such that
+The following theorem is a restatement of Theorem 2.1, with the notation described above.
+Theorem 6.5.
+Let where and let be the corresponding operator for defined as in (6.1). Then is locally bounded on for all and the family is uniformly continuous as
+Proof.
+Since we only seek local boundedness, we can assume that is small enough so that Assumption 1 holds.
+Let By the definition of Without loss of generality, we assume that is supported in the set for
+For every we have
+Since is compactly supported in the -variable, it follows that there is a distribution smooth in and supported at such that
+Since has homogeneity with respect to dilations in the -variable, we have for some Therefore, the integral kernel of satisfies
+where are defined as in Definition 5.9, and the sum converges pointwise. By Corollary 5.16, the sequence satisfies Conditions (I), (II) and (III) of Definition 3.1, with constants
+To verify the same for we need to use the fact that the pseudodifferential calculus is closed under adjoints.
+This is proved in [AMY22, Proposition 3.13], and moreover the adjoint is related to the -fibred distribution
+The corresponding cocycle is given by
+Since the function has all the same properties just proved for We have
+Here we are using Lemma 5.2 to replace by Therefore, also satisfies Conditions (I), (II) and (III) of Definition 3.1, with the same constants as
+Assume now that In this case, we have
+So for Corollary 5.16 implies that satisfies all of the conditions of Definition 3.1. Hence in this case, Corollary 4.3 and an easy limit argument proves that
+In general, for we can reduce to the case by replacing by using the fact that [AMY22, Proposition 3.14]. Thus in the general case we still have
+Since and are uniformly bounded as it follows from Corollary 4.2 that has weak-type uniformly as
+Hence, for general with and we have
+β
+Finally, we can conclude the part of Theorem 2.1
+Theorem 6.6.
+Let and let If then is locally bounded on
+Moreover, the associated family is uniformly continuous as in the sense that if then
+Proof.
+Theorem 6.5 is exactly this result stated in a coordinate chart Since the result is local, we immediately obtain the theorem. β
+7. Compactness of negative order operators
+The boundedness and compactness of negative order operators is much easier than zero order operators, since in this case we can represent the operator as a norm convergent series.
+Theorem 7.1.
+Let let for and let be the corresponding operator as in (6.1). Then is compact on and the family is uniformly continuous as
+Proof.
+Since the result is local, we may assume that satisfies Assumption 1. Since and are not included in Theorem 6.5, we first show the uniform boundedness of on as Let
+By definition and just as in Thoerem 6.5 for every we have
+Since is compactly supported in the -variable and has negative order, it follows that
+Therefore
+where is related to precisely as in Definition 5.9.
+By Lemma 5.12, the sequence satisfies Condition II.
+The adjoint is an operator of the same form, and hence satisfies the requirements of Lemma 4.4.
+Hence, by Lemma 4.4, we get
+Since we conclude the uniform boundedness of on for as
+Observe that the integral kernel of is smooth and compactly supported. In particular, is compact on Since is a norm convergent sum of compact operators, it follows that itself is compact. β
+We conclude with the part of Theorem 2.1.
+Theorem 7.2.
+Let and let where Then is locally bounded on and the associated family is uniformly continuous as on Moreover is locally compact.
+Proof.
+Since the assertion is local, it suffices to work in a single coordinate chart, and in this case the result is Theorem 7.1. β
+8. Sobolev spaces
+In this section we obtain results about the (local) boundedness of negative order operators between spaces, essentially we derive Sobolev embedding theorems for the AMY calculus.
+Definition 8.1.
+Let be a filtered manifold, and The local Sobolev space is the set of all distributions such that for any
+Example 8.2.
+Let be vector fields on which satisfy the rank HΓΆrmander property; that is, the tangent space of is everywhere spanned by the iterated brackets
+of depth at most Let be the filtration determined by etc.
+As in the previous sections, is an open set and are vector fields on equipped with weights We have defined
+We assume as before that satisfies Assumption 1.
+Lemma 8.3.
+Proof.
+The formula for is
+The family of functions has uniformly bounded -norm, hence over any compact set of there is such that
+β
+Corollary 8.4.
+For we have
+Proof.
+If is any filtration of vector fields on a manifold then near any point we can choose a weighted family of vector fields defining the filtration and hence a corresponding value of This integer is not really natural since a different choice of weighted vector fields could define the same filtration with a different value of However, for a given choice of we can derive a Sobolev embedding theorem for the AMY calculus.
+For this section we will assume that is a filtered manifold such that there exists a maximal so that near any point of there exists a weighted family of vector fields generating with sum of weights at most
+Lemma 8.5.
+Let and where Then maps locally compactly into
+Proof.
+Since the assertion is local, we may again work in a chart Write where as in Theorem 7.1. By Corollary 8.4, we have
+provided that Since each summand is smooth and compactly supported, is locally compact from to and the above computation shows that truncated to any compact subset of is a convergent sum of compact operators.
+β
+The following lemma represents a Hardy-Littlewood-Sobolev inequality for the filtration The scheme of the proof is essentially the same as classic interpolation-based proofs of Hardy-Littlewood-Sobolev, as in e.g.
+[Ste70, Theorem 1, Page 119].
+Lemma 8.6.
+Let where For all is locally bounded from into where
+Proof.
+It suffices to work in a single coordinate chart We show that is bounded from to the result will then follow by interpolation. Since we work locally, take
+Let We estimate
+Note that if then Markovβs inequality delivers
+So we concentrate on the case that Choose such that
+Let
+From the choice of and Corollary 8.4,
+Therefore
+By Markovβs inequality, it follows that
+By the choice of we have
+Hence
+That is, maps to and this completes the proof. β
+Theorem 8.7.
+Let be a filtered manifold with the property that there exists an integer so that near any point of there exists a weighted family of vector fields generating with sum of weights at most
+For all we have a locally compact embedding
+while for and
+we have a continuous embedding
+Proof.
+It was proved in [AMY22, Proposition 3.32] that for every there exists a pair of operators and such that By definition if then and therefore
+That is, is in the image of an operator of order β
+Example 8.8.
+Let be vector fields on an open set satisfying the HΓΆrmander condition of rank as in Example 8.2.
+Let be the graded dimension of the free nilpotent Lie algebra of depth with generators. Then for and supported on a compact subset of Theorem 8.7 and Example 8.2 imply that
+On the other hand, if we have
+In this example represents a worst-case-scenario and is in many cases much larger than necessary.
+9. Applications to elliptic regularity in
+Recall that the Sobolev space associated to a filtered manifold was given in Definition 8.1.
+The following is an immediate consequence of Theorem 2.1 and the composition properties of the AMY calculus.
+Lemma 9.1.
+If and If then is a continuous mapping
+One application of Lemma 9.1 in the unfiltered case is the a priori regularity theory for elliptic pseudodifferential operators, which can be proved via the fact that an elliptic operator of order has an inverse of order modulo smoothing operators (a parametrix).
+It was not proved in [AMY22] that a maximally hypoelliptic differential operator has a parametrix in the calculus, instead it was only proved for operators obeying the so-called strong -Rockland condition [AMY22, Definition 3.28].
+We do not need to recall the definition of that property here, we only record the following application of the existence of a parametrix.
+Theorem 9.2.
+Let let satisfy the strong -Rockland condition. We have the elliptic regularity
+Proof.
+Applications of elliptic regularity are well-known. As an example, the following is an existence and a priori regularity result for a certain non-linear PDE.
+Theorem 9.3.
+Let be a compact filtered manifold, equipped with a Riemannian metric and let be a symmetric operator with respect to the inner product on induced by Assume that satisfies the strong -Rockland condition.
+Let For all real-valued there exists such that